Calculate the molality NH3 (17 g/mol) in a 5.05% (w/w) aqueous solution. (dsoln = 0.9651 g/mL) three sig figs
Answer #1

5.05 w/w % means

5.05 g NH3 present in 100 g solution

mass of solute = 5.05

mass of solvent = 100 - 5.05 = 94.95 g

molality = (W /MW) (1000 / mass of solvent in g)

molality = (5.05 / 17) (1000 / 94.95)

molality = 3.13 m